please help

how much water must be mixed with 8 pints of a solution that is 90% developer to produce a mixture that is 26% developer?

Answer by checkley79(3341)

.9*8+X=.26(8+X)
.72+X=2.08+.26X
.72+X-.26X=2.08
.74X=2.08-.72
.74X=1.36
X=1.36/.74
X=1.84 PINTS OF WATER ARE ADDED.
PROOF:
.9*8+1.84=.26(8+1.84)
.72+1.84=.26*9.84
2.56=2.56